OpenStudy (anonymous):

use rational exponent to simplify 5sqrtx^2=x^4?

OpenStudy (anonymous):



OpenStudy (anonymous):

x^20? or x^2?

OpenStudy (anonymous):

yes its x^20

OpenStudy (anonymous):

ok so is it the fifth root of x^20? like this \[\sqrt[5]{x^{20}}=x^4\] or like this? \[5\sqrt{x^{20}}=x^4\]